4. You need to borrow $7,500 to buy a motorcycle. The dealer offers an APR of 9.25% to be paid off in monthly installments over 3 years.
(a) What is the monthly payment? (7)
(b) How much total interest did you pay? (7)
Answer #1

We need to calculate EMI for the above. We can do it two ways:

1. Manually:

We compute PV factor for n = 3 Years * 12 months => 36, and Rate r = 9.25% / 12 = 0.7708 %. We can use a normal calculator since we may not get the PV factors for the given rate.  So, we get 0.7708% for 36 = 31.332.

Accordingly, We pay $ 7500 at end of 36 instalments. Therefore EMI => 7500 / 31.332 = $ 239.37.

2. Using excel, We use formula =PMT(Rate, Nper, Principal)

ie., =PMT( 0.0925/12, 3*12, 7500) and we get EMI = $ 239.37

a. Monthly Payment = $ 239.37.

b. Total Interest =>

Total EMIs (239.37*36) 8617.32
Less: Principal 7500.00
Total Interest 1117.32